What is the z score for each number: 50,75,75,100
Answer by edjones(8007)

M=(50+75+75+100)/4=75
v=((50-75)^2+(75-75)^2*2+(100-75)^2)/3 = 1250/3
sqrt(v)=s=20.41
z=(x-m)/s
z[50]=-25/20.41 = -1.22
z[75]=0
z[100]=25/20.41 = 1.22
.
Ed